# Dynamic Contextual Pricing with Doubly Non-Parametric Random Utility Models

### v1

@@ 4.2 Real Data Analysis
We utilize the real-life auto loan dataset from Columbia University’s Center for Pricing and Revenue Management, a dataset also employed in several prior studies (Phillips et al., 2015; Ban and Keskin, 2021; Luo et al., 2023; Wang et al., 2023). This dataset encompasses 208,085 auto loan applications spanning from July 2002 to November 2004 and includes various features, such as loan amount and borrower information. For consistency, we perform feature selection in line with (Ban and Keskin, 2021; Luo et al., 2023; Wang et al., 2023), focusing on the following four features: approved loan amount, FICO score, prime rate, and competitor’s rate. As for the price variable, we also computed it in the same way as the aforementioned literature, where $p_{t}=\textrm{Monthly Payment}\cdot\sum_{t=1}^{\textrm{Term}}(1+\textrm{Rate})^{-t}-\textrm{Loan Amount}$. The rate is set as $0.12\%$, which is an approximate average of the monthly London interbank rate for the studied time period. Moreover, this dataset also records the purchasing decisions of the borrowers, given the price set by the lender.
@@
Note that one is not able to obtain online responses to any algorithms. Thus, we follow the calibration idea proposed in (Ban and Keskin, 2021; Luo et al., 2023; Wang et al., 2023) to learn a binary choice model using the entire dataset and leverage it as the ground truth to conduct online numerical experiments. To make calibration model more stable, we first process all covariates by standardization and take log of the price variable, one can also remove the extreme values that are outside the 0.01 and 0.99 percentiles to mitigate the influence of outliers. For the learning decision model, we typically fit a generalized logistic regression with a quadratic form as the logits; meanwhile, instead of using a conventional logistic function, we adopt a specified sigmoid function with bounded support. To be specific, we denote the contextual covariates as $\{\boldsymbol{x}_{t}=(x_{t1},x_{t2},x_{t3},x_{t4})\}_{t\geq 1}$, and the mean utility function of form $\mathring{\mu}(\boldsymbol{x}_{t})=\beta_{1}x_{t1}^{2}+\beta_{2}x_{t2}^{2}+\beta_{3}x_{t3}^{2}+\beta_{4}x_{t4}^{2}$.
@@
With the belief that the coefficients of the mean utility function can be varied across the data while the marketing noise shares the same distribution. We employ a two-stage approach learning calibration model. Initially, the covariates $\{\boldsymbol{x}_{t}\}_{t\geq 1}$ are divided into 20 clusters based on similarities, utilizing the K-means clustering technique. Subsequently, the model applies a generalized logistic regression to each cluster individually to learn the coefficients of the quadratic form within each group. Note that the marketing noise $\varepsilon_{t}$ is set to follow the same distribution with probability density function $f(v)\propto(1-(\frac{v}{3})^{2})^{4}\cdot\mathbbm{1}(\left\lvert v\right\rvert\leq 3).$ Figure 3 demonstrates the alignment between observed data and the theoretical model over the marketing noise $\varepsilon_{t}$, which means the calibration model is well specified.
@@
Given these key components, the remaining experiments remain almost the same as discussed in the simulation part. Here we use the calibration model, setting $\mathring{\mu}(\boldsymbol{x}_{t})$ with parameter $\boldsymbol{\beta}_{k},k=1,\dots,20$, and the noise distribution $f(v)$ given above and sample $\boldsymbol{x}_{t}$ from those four features. Since in reality we don’t know the true smoothness order of noise distribution, in each episode, we set the exploration phase with length $|{\cal T}_{k,exp}|=\lfloor(dn_{k})^{\frac{d+4}{d+8}}\rfloor$. We set $B=6,n_{0}=200$ and conduct Algorithm 1. Recall from Theorem 7 that the regret $\operatorname{Reg}(\pi,T)\lesssim[(\log T)^{2}+d\log T]T^{\frac{d+4}{d+8}}.$ Hence, we plot $\widetilde{\textrm{reg}}(T)$ against $\log(T)-\log(1500)$ in Figure 4(b), where $\widetilde{\textrm{reg}}(T):=\log(\operatorname{Reg}(T))-(2\log\log T+\log 4\log T)-(\log(\operatorname{Reg}(T=1500))-2\log\log 1500-\log 4\log 1500)$.
Figure 4: Comparison between our methods and linear model based policy and ‘RMLP-2’ based on real data application. (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. The dashed lines has slope equal to $\frac{d+4}{d+8}$ with $d=4$. Here we only run algorithm one time.

### v2

@@ 4.2 Real Data Analysis
We utilize the real-life auto loan dataset from Columbia University’s Center for Pricing and Revenue Management, a dataset that has also been employed in several prior studies (Phillips et al., 2015; Ban and Keskin, 2021; Luo et al., 2024; Wang et al., 2023). This dataset encompasses 208,085 auto loan applications spanning from July 2002 to November 2004 and includes various features, such as loan amount and borrower information. For consistency, we perform feature selection in line with existing work (Ban and Keskin, 2021; Luo et al., 2024; Wang et al., 2023), focusing on the following four features: approved loan amount, FICO score, prime rate, and competitor’s rate. Regarding the price variable, we also computed it in the same way as the aforementioned literature, where $p_{t}=\textrm{Monthly Payment}\cdot\sum_{t=1}^{\textrm{Term}}(1+\textrm{Rate})^{-t}-\textrm{Loan Amount}$. The rate is set as $0.12\%$, which is an approximate average of the monthly London interbank rate for the studied time period. Moreover, this dataset also records the purchasing decisions of the borrowers, given the price set by the lender.
@@
Note that one is not able to obtain online responses to any algorithms. Thus, we follow the calibration idea proposed in (Ban and Keskin, 2021; Luo et al., 2024; Wang et al., 2023) to first learn a binary choice model using the entire dataset and then leverage it as the ground truth to conduct numerical experiments. To make calibration model more stable, we process all covariates by standardization and take log of the price variable. One can also remove the extreme values that are outside the 0.01 and 0.99 percentiles to mitigate the influence of outliers. For learning the decision model, we typically fit a generalized logistic model with a quadratic form as the logits; meanwhile, instead of using a conventional logistic function, we adopt a specified sigmoid function with bounded support.
@@
Specifically, denote the contextual covariate as $\boldsymbol{x}_{t}\in\mathbb{R}^{4},\;t=1,2,\dots$ and the quadratic form $\mathring{\mu}(\boldsymbol{x}_{t})=\boldsymbol{\beta}^{\top}\boldsymbol{x}_{t}^{2},\;\boldsymbol{\beta}\in\mathbb{R}^{4}$ for the mean utility function. Then the noise $\{\varepsilon_{t}\}$ are set to independently follow the distribution with probability density function $f(z)\propto(1-(\frac{z}{3})^{2})^{4}\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 3).$ Furthermore, with the belief that the coefficients of the mean utility function can be varied across the data while the noise shares the same distribution. We employ a two-stage approach learning calibration model. Initially, the covariates $\{\boldsymbol{x}_{t}\}_{t\geq 1}$ are divided into 20 clusters based on similarities, utilizing the K-means clustering technique. Here we denote $\boldsymbol{x}_{t}^{(k)}$ if the covariate $\boldsymbol{x}_{t}$ belongs to the $k$-th cluster. Then, we apply generalized logistic regression to each data cluster, fitting the coefficients of the quadratic form for each group while maintaining the same noise distribution. Figure 3 demonstrates the alignment between observed data and the theoretical model over the noise $\varepsilon_{t}$, which means the calibration model is well specified.
@@
Given these key components, the remaining experiments remain almost the same as discussed in the simulation part. Here the calibration model adopts the form $\mathring{\mu}(\boldsymbol{x}_{t})=\mathring{\mu}(\boldsymbol{x}_{t}^{(k)})=\boldsymbol{\beta}_{k}^{\top}\big{(}\boldsymbol{x}_{t}^{(k)}\big{)}^{2},\;k=1,\dots,20$, where $\boldsymbol{\beta}_{k}$ is the learned coefficients when the contextual covariate $\boldsymbol{x}_{t}$ belongs to the $k$-th cluster. Since in reality we don’t know the true smoothness order of noise distribution, in each episode, we set the exploration phase with length $|{\cal T}_{k,exp}|=\lfloor(dn_{k})^{\frac{d+4}{d+8}}\rfloor$. We set $B=6,n_{0}=200$ and conduct Algorithm 1. Recall from Theorem 8 that the regret $\operatorname{Reg}(\pi,T)\lesssim[(\log T)^{2}+d\log T]T^{\frac{d+4}{d+8}}.$ Hence, we plot $\widetilde{\textrm{reg}}(T)$ against $\log(T)-\log(1500)$ in Figure 4(b), where $\widetilde{\textrm{reg}}(T):=\log(\operatorname{Reg}(T))-(2\log\log T+\log 4\log T)-(\log(\operatorname{Reg}(T=1500))-2\log\log 1500-\log 4\log 1500)$.
Figure 4: Comparison between our methods and linear model based policy and ‘RMLP-2’ based on real data application. (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. The dashed lines has slope equal to $\frac{d+4}{d+8}$ with $d=4$.